OpenStudy (mishapocolypse0314):

WILL FAN AND MEDAL!!! evaluate; sin(2cos^-1(1/sqrt5))

OpenStudy (anonymous):

u still need help

OpenStudy (mishapocolypse0314):

yes please

OpenStudy (anonymous):

ok let me try to figure it out

OpenStudy (anonymous):

an(t) = y/x = (2sqrt(6))/1 = 2sqrt(6)

OpenStudy (zarkon):

use \[\sin(2\theta)=2\sin(\theta)\cos(\theta)\]

OpenStudy (mishapocolypse0314):

thank you soooo much!!!

OpenStudy (anonymous):

what no fan and medal for me lol

OpenStudy (mishapocolypse0314):

sorr bro but the back of the book said the answer was 4/5 lol

OpenStudy (anonymous):

wow for real

OpenStudy (zarkon):

Just so you know ... all mods can see deleted posts

OpenStudy (anonymous):

boom deleted

OpenStudy (anonymous):

why erase

OpenStudy (mishapocolypse0314):

then why the hell dont you get off and do your homework on your own

OpenStudy (anonymous):

sleek-feathered one im try in to help u scum retricefeatheres like u sleek-feathered one

OpenStudy (anonymous):

u dirk

OpenStudy (anonymous):

so shut the hell up

OpenStudy (anonymous):

ban me u retricewhipes i dont need u

OpenStudy (mishapocolypse0314):

im not the one getting butthurt over a medal

OpenStudy (anonymous):

u bitc1h i dont need a medal the hell with medals they dont get u knowthing tom cruis

OpenStudy (anonymous):

get a job u bum and get the hell of openstudy its going down in the bm bitc1hes

OpenStudy (anonymous):

come on u chicken say something

OpenStudy (anonymous):

mane just forget it im block in u and im out